OpenStudy (anonymous):

Solve the systems using substitution. show me how please 6x + 2y = 0 4x + y = -1

OpenStudy (anonymous):

can you solve the top one for y

OpenStudy (anonymous):

no i dont know how to do any of this ''/

OpenStudy (anonymous):

okay ill go through the steps first: 6x + 2y = 0 2y = -6x (i subtracted 6x from both sides) y = -3x (i divided both sides by 2)

OpenStudy (anonymous):

okay then whats next?

OpenStudy (anonymous):

now you know y = -3x so you can plug this into the second equation: 4x + y = -1 and get 4x + (-3x) = -1 x = -1 (just added 4x and -3x)

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

kind of like i get the steps but the equation left makes no sense lol is it suppose to?

OpenStudy (anonymous):

well when solveing these two equations your trying to find an x and y that make both true at the same time. what weve got is that the x value that does this is -1.

OpenStudy (anonymous):

the y value can be found by plussing in x = -1 into either of the equations and solving for y

OpenStudy (anonymous):

can you do that?

OpenStudy (anonymous):

so confusing :( see my teacher gives me a web site to put the problem in and it gives me he answer he tells me o cheat but i dont want to, so sorry if it annoys you but i still dont get it :(

OpenStudy (anonymous):

thats no problem. do you understand up to the point of where we got x = -1

OpenStudy (anonymous):

if not its ok

OpenStudy (anonymous):

no cus 4 +-3 =1 not -1...... ''/

OpenStudy (anonymous):

im trying to learn because i have 5 more problems like this to solve before the night is up and i have to show work

OpenStudy (anonymous):



OpenStudy (anonymous):

okay thats better im going to try and solve the rest if i have any problems can i come to you ?

OpenStudy (anonymous):

yeah, just tag me in this post or send me a message. dont know how late ill be up but i will help you while i can : )